Ball a and b are 20m from ground. Ball a was thrown upwards with a speed of 5m/s. Ball b was thrown at -4m/s. find the maximum height
1
Expert's answer
2019-09-10T13:43:30-0400

The maximum height will have the ball a:


"H_a=h+\frac{v_a^2}{2g}"

"H_a=20+\frac{5^2}{2(9.8)}=21.3\ m"